# Less is More: Local Intrinsic Dimensions of Contextual Language Models

### v1

@@ A.2 Noise Analysis
(a) Global TwoNN estimate.
(b) Mean of local TwoNN estimates.
(c) Standard deviation of local TwoNN estimates.
Figure 9: Effect of artificial Gaussian noise on masked token embeddings. On the $x$-axis we mark the Hausdorff distance of the noisy embeddings to the clean embeddings, with color denoting the distortion $\sigma$ in the artificial Gaussian noise.
@@
[…]
Gaussian noise is applied to all dimensions of the embeddings with varying distortion parameters, $\sigma\in\{0.001;0.002;0.003;0.004;0.01\}$, and multiple random seeds to introduce variability. For each noise level, we compute the approximate Hausdorff distance between the point cloud of the 60 0006000060\,00060 000 clean embedding vectors and the 60 0006000060\,00060 000 noisy embedding vectors, using the Euclidean distance as the distance metric.
The results are visualized in Figure 9, where we plot the Hausdorff distance against three measures: the global TwoNN-estimate of the noisy point cloud, the mean of the local TwoNN estimates, and the standard deviation of the local TwoNN estimates.
@@ A.3 Masked Token Embeddings versus Regular Token Embeddings
The plots in Figure 10 compare the TwoNN estimates for regular tokens (left) and masked tokens (right) in the MultiWOZ dataset at the last layer. Each box-plot represents the distribution of token-wise TwoNN estimates.
@@ Appendix B Additional details on datasets used
See Table 1 for details on the dataset cards and corresponding split sizes.
@@
MultiWOZ2.1 https://huggingface.co/datasets/ConvLab/multiwoz21 113 556113556113\,556113 556 14 7481474814\,74814 748 14 7441474414\,74414 744
EmoWOZ https://paperswithcode.com/dataset/emowoz-1 66 4746647466\,47466 474 $8509$ $8634$
SGD https://huggingface.co/datasets/google-research-datasets/schema_guided_dstc8 329 964329964329\,964329 964 48 7264872648\,72648 726 84 5948459484\,59484 594
Wikipedia (random 80%-10%-10% split) https://huggingface.co/datasets/Salesforce/wikitext/viewer/wikitext-103-v1 1 441 08014410801\,441\,0801 441 080 180 135180135180\,135180 135 180 135180135180\,135180 135
Reddit (random 80%-10%-10% split) https://huggingface.co/datasets/SocialGrep/one-year-of-tsla-on-reddit 174 996174996174\,996174 996 21 8742187421\,87421 874 21 8742187421\,87421 874
@@ C.1 Additional fine-tuning results
Continuing the discussion in Section 4.1, in Figures 11, 12 and 13, we provide additional distribution plots of the local estimates. Figure 11 is interesting because here the RoBERTa model is fine-tuned on data that it has not seen during pre-training.
@@
Moreover, in Figures 12 and 13 we show the distribution of local estimates for the auto-regressive GPT-2-medium model in a natural language modeling task, and their shift under fine-tuning. The GPT-2-medium fine-tuning here is performed with standard parameters on the first 10 0001000010\,00010 000 sequences of the given dataset’s train portion, and we show the local estimates at the checkpoint after $1200$ batches.
(a) TwoNN estimates on MultiWOZ validation embeddings.
(b) TwoNN estimates on SGD validation embeddings.
(c) TwoNN estimates on Reddit validation embeddings.
Figure 11: Comparison of the intrinsic dimension (ID) across three data modalities for both models. The ID of embeddings originating from the fine-tuning distribution (Reddit) differs markedly between models, whereas the IDs for the out-of-distribution corpora (SGD, MultiWOZ) are almost indistinguishable.
(a) TwoNN estimates on MultiWOZ validation embeddings.
(b) TwoNN estimates on Wikipedia validation embeddings.
(c) TwoNN estimates on Reddit validation embeddings.
Figure 12: Comparison of the intrinsic dimension (ID) across three data modalities for both models. The ID of embeddings originating from the fine-tuning distribution (MultiWOZ) differs markedly between models, whereas the IDs for the out-of-distribution corpora (Wikipedia, Reddit) are almost indistinguishable.
(a) TwoNN estimates on MultiWOZ validation embeddings.
(b) TwoNN estimates on Wikipedia validation embeddings.
(c) TwoNN estimates on Reddit validation embeddings.
Figure 13: Comparison of the intrinsic dimension (ID) across three data modalities for both models. The ID of embeddings originating from the fine-tuning distribution (Wikipedia) is markedly lower, whereas the IDs for the out-of-distribution corpora (MultiWOZ, Reddit) are even higher for the fine-tuned variants.
C.2 Layerwise computation of local estimates
@@
Our setup from Section 3 and algorithm 1 naturally applies to embeddings derived from arbitrary layers of the language model. In particular, the mean local estimates can be compared layer-wise between different checkpoints of a model, as we show in Figure 14. The main observations here are that the drop in mean local dimension on the dataset used for fine-tuning is visible over all the model layers. It appears to be most pronounced in the intermediate and last layers of the model.
@@
1212-12- 121111-11- 111010-10- 1099-9- 988-8- 877-7- 766-6- 655-5- 544-4- 433-3- 322-2- 211-1- 15555101010101515151520202020LayerMeanTwoNN dimension estimates for MultiWOZ-validationRoBERTaRoBERTa fine-tuned on MultiWOZ-train1212-12- 121111-11- 111010-10- 1099-9- 988-8- 877-7- 766-6- 655-5- 544-4- 433-3- 322-2- 211-1- 15555101010101515151520202020LayerMeanTwoNN dimension estimates for Wikipedia-validationRoBERTaRoBERTa fine-tuned on MultiWOZ-train
@@
Figure 14: Comparison of the development of the ID through the different layers of the two models. The ID of embeddings from the distribution that has been used for fine-tuning differs significantly between the two models, whereas the ID of other data distributions are indistinguishable.
@@ C.3 Additional Grokking results
This section supplements the grokking experiments in Section 4.2. See Figure 15 for a plot comparing model performance and mean local estimates on the training set, shown separately for the training run seeds. This representation is useful, as the timing of the onset of grokking can be very dependent on the model and dataset shuffling seed. It remains true that for individual seeds, the drop in mean local dimension on the training set coincides with the increase in validation accuracy.
@@
Figure 15: Training a model on addition mod $p=197$ with different training data fraction selected from $\{0.1;0.15;0.2;0.25;0.3;0.4;0.5\}$. Five seeds per configuration are shown separately, demonstrating that the onset of grokking and the drop in mean local dimension vary greatly depending on the specific training run. The plots show the development over 15 0001500015\,00015 000 batches. The mean local estimates are computed on the training split for $N=3000$; $L=64$.
@@ C.4 Additional Trippy-R dialogue state tracking results
We here show additional results for Section 4.3, where we discuss exhausting model training capabilities. See Figure 16(c) for longer TripPy-R dialog state tracking runs over 50 epochs. After about $20$ training epochs, the mean local estimates stabilize and predict the convergence of the model performance on the downstream task.
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The paper has four main claims, which are listed in the last paragraph of the Introduction Section 1. Each claim corresponds to a subsection of the Experiments Section 4.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: In the methods section, we mention limitations whenever they lead to restrictions in our setup, for instance, in Section 3.4. We additionally dedicate Section 4.6 to a separate discussion of limitations.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We do not provide new theoretical results in this work. For those local dimension estimation methods that we apply, in Section 3.4, we clearly state references to the original publications that contain a detailed discussion of the assumptions.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Our method is described in full detail in Section 3 and in algorithm 1. The experiment sections Sections 4.1, 4.2, 4.3 and 4.4 contain all the details necessary to reproduce the model training runs, which form the basis of our main findings. Moreover, our code will be released as part of the supplemental material.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Our code is submitted as part of the supplemental material and will be made public upon publication. All datasets are freely available under permissive licenses; more details can be found in the experiments section Section 4 and Appendix B.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Training and test details, including hyperparameters and details about splits are described in the experiment setup sections Sections 4.1, 4.2, 4.3 and 4.4.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: To confirm the findings of our main results, we run every experimental setup for several seeds, and report confidence intervals or standard deviation in Figures 2, 3 and 4. All standard deviations in this paper are sample standard deviations, computed with the numpy and pandas library functions np.std(ddof=1) and pd.std(). Our sensitivity analysis in Appendix A studies the effect of sampling size and locality parameters by reporting the results for multiple sub-sampling seeds per hyperparameter choice. The violin plots of the local estimates in Figure 1 and Section C.1 show the entire distribution and are reported together with mean and standard deviations.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We discuss the runtime and computational resources in Section 4.5.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We reviewed the NeurIPS Code of Ethics and confirm that we have followed every aspect.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: Our work is foundational and provides general insights into the training dynamics of language models. While we acknowledge the general risks involved in developing and deploying machine learning models, our insights do not lead to new negative societal implications. Since we aim to geometrically understand the internal mechanism in a language model under different fine-tuning and training regimes, we hope that our work can help in safety and interpretability research in the future.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We believe our foundational work does not pose such safety risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite the original research papers introducing the models used in Sections 3.1, 4.2, 4.3 and 4.4 and acknowledge code that we used as the basis for our work in the submitted supplemental material. All datasets discussed are freely available under permissive licenses, and we provide links to the dataset cards in Appendix B.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Our code, which is submitted as part of the supplemental material and will be made public upon publication, is documented in the repository. This includes example hydra configuration files for the datasets and model selection. The Python package management system uv is used to reproduce the virtual environment with all dependencies, and run commands for the most important entry points into the codebase are provided.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We employ LLMs as standard components of the machine learning models under consideration, and describe their usage in detail in Sections 3 and 4. Our development process did not involve LLMs in a non-standard way.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@ A.2 Noise Analysis
(a) Global TwoNN estimate
(b) Mean of local TwoNN estimates
(c) Standard deviation of local TwoNN estimates
Figure 9: Effect of artificial Gaussian noise on masked token embeddings. On the $x$-axis, we mark the Hausdorff distance of the noisy embeddings to the clean embeddings, with color denoting the distortion $\sigma$ in the artificial Gaussian noise.
@@
We analyze the effect of Gaussian noise on the embeddings from the MultiWOZ test set using a fixed set of parameters and varying noise distortions. Specifically, we use M=10  [^0] M=\$10\,000\$ random sequences from the MultiWOZ test set and extract the masked token embeddings from the last layer of the RoBERTa base model, which produces 768-dimensional vectors. To compute local estimates, we sub-sample N=60  [^0] N=\$60\,000\$ points from the embeddings and set the neighborhood size to $L=128$.
Gaussian noise is applied to all dimensions of the embeddings with varying distortion parameters, $\sigma\in\{0.001;0.002;0.003;0.004;0.01\}$, and multiple random seeds to introduce variability. For each noise level, we compute the approximate Hausdorff distance between the point cloud of the 60 00060\,000 clean embedding vectors and the 60 00060\,000 noisy embedding vectors, using the Euclidean distance as the distance metric.
The results are visualized in Figure˜9, where we plot the Hausdorff distance against three measures: the global TwoNN-estimate of the noisy point cloud, the mean of the local TwoNN estimates, and the standard deviation of the local TwoNN estimates.
@@ A.3 Masked Token Embeddings versus Regular Token Embeddings
The plots in Figure˜10 compare the TwoNN estimates for regular tokens (left) and masked tokens (right) in the MultiWOZ dataset at the last layer. Each box plot represents the distribution of token-wise TwoNN estimates.
@@ Appendix B Additional Details on Datasets Used
See Table˜1 for details on the dataset cards and corresponding split sizes.
@@
MultiWOZ2.1 https://huggingface.co/datasets/ConvLab/multiwoz21 113 556113\,556 14 74814\,748 14 74414\,744
EmoWOZ https://paperswithcode.com/dataset/emowoz-1 66 47466\,474 $8509$ $8634$
SGD https://huggingface.co/datasets/google-research-datasets/schema_guided_dstc8 329 964329\,964 48 72648\,726 84 59484\,594
ICLR 2024 submissions (titles + abstracts) $5764$ $721$ $720$
Reddit (random 80%-10%-10% split) https://huggingface.co/datasets/SocialGrep/one-year-of-tsla-on-reddit 174 996174\,996 21 87421\,874 21 87421\,874
Wikipedia (random 80%-10%-10% split) https://huggingface.co/datasets/Salesforce/wikitext/viewer/wikitext-103-v1 1 441 0801\,441\,080 180 135180\,135 180 135180\,135
@@ C.1 Additional Fine-Tuning Results
Continuing the discussion in Section˜4.1, in Figure˜11, we provide additional distribution plots of the local estimates for fine-tuned RoBERTa-base models. Here and in the following, the plots labeled with FT correspond to embedding spaces created from the respective base model fine-tuned on the given dataset (MultiWOZ, Reddit, Wikipedia).
(a) TwoNN estimates on MultiWOZ Validation embeddings
(b) TwoNN estimates on SGD Validation embeddings
(c) TwoNN estimates on Reddit Validation embeddings
(d) TwoNN estimates on Wikipedia Validation embeddings
Figure 11: Distribution of LIDs of the masked language model RoBERTa-base and fine-tunes (FT).
@@
Moreover, in Figure˜12 we show the distribution of local estimates for the auto-regressive GPT-2-medium model, and their shift under fine-tuning on a causal language modeling task. The GPT-2-medium fine-tuning here is performed with standard parameters on the first 10 00010\,000 sequences of the given dataset’s train portion, and we show the local estimates at the checkpoint after $1200$ batches.
Interesting cases to note are the RoBERTa and GPT-2-medium models fine-tuned on the Reddit data: Since the text from the dataset was created in the year 2022, long after the training of RoBERTa and GPT-2-medium finished, this combination illustrates the behavior of a model fine-tuned on data that it has not seen during pre-training. The general observations from Section˜4.1 hold for the autoregressive models as well: The LIDs of embeddings originating from the fine‑tuning distribution decrease markedly between models, whereas the LIDs for the out‑of‑distribution corpora remain largely unaffected or have increased.
(a) TwoNN estimates on MultiWOZ Validation embeddings
(b) TwoNN estimates on SGD Validation embeddings
(c) TwoNN estimates on ICLR 2024 Validation embeddings
(d) TwoNN estimates on Reddit Validation embeddings
(e) TwoNN estimates on Wikipedia Validation embeddings
Figure 12: Distribution of LIDs of the autoregressive GPT-2-medium model and fine-tunes (FT).
Local estimates for larger models
The geometric perspective and methodology we propose are broadly applicable across model architectures. Local estimates computations for the Phi-3.5-mini-instruct model in Figure˜13 and for Llama-3.1-8B in Figure˜14 demonstrate that our sub-sampling procedure yields meaningful estimates even for latent spaces whose ambient dimension is in the thousands, a typical setting for modern LLM architectures.
Phi-3.5-mini-instruct is a decoder model with 3.82B parameters and a hidden dimension of $3072$ (Abdin et al.,, 2024), while Llama-3.1-8B is even larger with 8.03B parameters and a hidden dimension of $4096$ (AI@Meta,, 2024). We fine-tune via LoRA with rank constraint $r=16$ on a random subsample of 10 00010\,000 sequences from the MultiWOZ and Reddit training splits, with other hyperparameters as described in Section˜4.1. The violin plots in Figures˜13 and 14 show the distribution of the LIDs of the validation split for the model checkpoint after 800 batches.
(a) TwoNN estimates on MultiWOZ Validation embeddings
(b) TwoNN estimates on SGD Validation embeddings
(c) TwoNN estimates on ICLR 2024 Validation embeddings
(d) TwoNN estimates on Reddit Validation embeddings
(e) TwoNN estimates on Wikipedia Validation embeddings
Figure 13: Distribution of LIDs of the autoregressive Phi-3.5-mini-instruct model and fine-tunes (FT).
(a) TwoNN estimates on MultiWOZ Validation embeddings
(b) TwoNN estimates on SGD Validation embeddings
(c) TwoNN estimates on ICLR 2024 Validation embeddings
(d) TwoNN estimates on Reddit Validation embeddings
(e) TwoNN estimates on Wikipedia Validation embeddings
Figure 14: Distribution of LIDs of the autoregressive Llama-3.1-8B model and fine-tunes (FT).
C.2 Layer-Wise Computation of Local Estimates
@@
Our setup from Section˜3 and Algorithm 1 naturally applies to embeddings derived from arbitrary layers of the language model. In particular, the mean local estimates can be compared layer-wise between different checkpoints of a model, as we show in Figure˜15. The main observations here are that the drop in mean local dimension on the dataset used for fine-tuning is visible over all the model layers. It appears to be most pronounced in the intermediate and last layers of the model.
@@
12-1211-1110-109-98-87-76-65-54-43-32-21-155101015152020LayerMeanTwoNN dimension estimates for MultiWOZ-validationRoBERTaRoBERTa fine-tuned on MultiWOZ-train12-1211-1110-109-98-87-76-65-54-43-32-21-155101015152020LayerMeanTwoNN dimension estimates for Wikipedia-validationRoBERTaRoBERTa fine-tuned on MultiWOZ-train
@@
Figure 15: Comparison of the mean LID development through the different layers of RoBERTa-base and a fine-tuned variant. The LID of embeddings from the distribution used for fine-tuning (here MultiWOZ) differs significantly between the two models, whereas the LID of other data distributions (here Wikipedia) is indistinguishable.
@@ C.3 Additional Grokking Results
This section supplements the grokking experiments in Section˜4.2. See Figure˜16 for a plot comparing model performance and mean local estimates on the training set, shown separately for the training run seeds. This representation is useful, as the timing of the onset of grokking can be very dependent on the model and dataset shuffling seed. It remains true that for individual seeds, the drop in mean local dimension on the training set coincides with the increase in validation accuracy.
@@
Figure 16: Training a model on addition mod $p=197$ with different training data fraction selected from $\{0.1;0.15;0.2;0.25;0.3;0.4;0.5\}$. Five seeds per configuration are shown separately, demonstrating that the onset of grokking and the drop in mean local dimension vary greatly depending on the specific training run. The plots show the development over 15 00015\,000 batches. The mean local estimates are computed on the training split for $N=3000$; $L=64$.
@@ C.4 Additional TripPy-R Dialogue State Tracking Results
We here show additional results for Section˜4.3, where we discuss exhausting model training capabilities. See Figure˜17(c) for longer TripPy-R dialog state tracking runs over 50 epochs. After about $20$ training epochs, the mean local estimates stabilize and suggest the convergence of the model performance on the downstream task.